a rectangle has an area of 24 square units. the width is 5 units less than the length. what is the lenght in units of the rectangle?
Found 3 solutions by hokies, palanisamy, Alan3354:
Answer by hokies(65)

L * W = 24
W = L - 5
Solve this system of equations (plug the 2nd equation into the 1st equation): L=8, W=3

Answer by palanisamy(496)

Let the length of the rectangle be x
Given, the width is 5 units less than the length.
So the breadth of the rectangle = x-5
Area of the rectangle is x(x-5) = 24
x^2-5x-24 = 0
(x-8)(x+3) = 0
x-8 = 0 or x+3 = 0
x = 8 or x = -3
x cannot be negative
So we get x = 8
Therefore the length of the rectangle = 8 miles
and the breadth of the rectangle = x-5 = 8-5 = 3 miles

Answer by Alan3354(30993)

a rectangle has an area of 24 square units. the width is 5 units less than the length. what is the lenght in units of the rectangle?
-------------------
L = 8
W = 3